OpenStudy (anonymous):

a piece of wire 20 meters long is cut into 2 pieces. one piece is bent into the shape of a square. the other is bent in the shape of an equilateral triangle. what length of wire should be used to form the square in order to ensure that the total are enclosed within the 2 shapes is in a minimum?? I am so confused with this!

OpenStudy (anonymous):

total area?

OpenStudy (anonymous):

is in a minimum or is a minimum

OpenStudy (anonymous):

initial length =L piece for a square length = S piece for a triangle length = T Total are enclosed = A (1/4)S*(1/4)S + sqrt(3)/2T*T = A S+T = L T=L-S so: (1/4)S*(1/4)S + sqrt(3)/2(L-S)^2 = A A will have a minimum at the point where A' =0 so find the derivative and set it to 0, to find S at which it happens.

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

it doesnt give me the total area, just the minimum. and what does the minimum mean?

OpenStudy (anonymous):